OpenStudy (hba):

If a*b equals the remainder when a+b is divided by 12,then (11*10)*9=................

OpenStudy (hba):

@sirm3d yes ?

OpenStudy (anonymous):

IS 6 THE ANSWER

OpenStudy (sirm3d):

yes it is.

OpenStudy (hba):

Yes but how do i do it lol :D

OpenStudy (sirm3d):

(11 * 10) * 9 = (9) * 9 = 6

OpenStudy (sirm3d):

just perform the operation inside the parentheses first

OpenStudy (hba):

please dont confuse me

OpenStudy (sirm3d):

what is the value of 11 * 10 ?

OpenStudy (hba):

110

OpenStudy (hba):

or 11+10/2

OpenStudy (sirm3d):

that is not multiplication

OpenStudy (anonymous):

its simple firstly solve the bracket 110 now divide it by 12 remainder is 2 multiply 2 by 9,its 18 divide 18 by 12 remainder is 6

OpenStudy (sirm3d):

a * b means you add a and b (a +b), divide by 12, and get the remainder

OpenStudy (sirm3d):

@nitz a * b = the remainder when (a + b) is divided by 12.

OpenStudy (hba):

SHOW ME THE WHOLE SOLUTION 0-0

OpenStudy (sirm3d):

11 * 10 means add the two numbers, divide the sum by 12, and get the remainder. so 11 * 10 = remainder in (11+10) divided by 12 or 11 * 10 = 2

OpenStudy (freckles):

Let a*b=R So that means we have \[\frac{a+b}{12}= Q+\frac{R}{12}\] Q is for Quotient I'm about to replace the remainder, R with a*b So we have \[12 \frac{a+b}{12}=12(Q+\frac{a*b}{12})\] \[a+b=12Q+a*b\] Solving for a*b gives us: \[a*b=a+b-12Q\]

OpenStudy (freckles):

\[11 * 10=11+10-12Q\] \[11*10=21-12Q\] Q=1 so we don't get a negative \[21-12(1)=21-12=9\]

OpenStudy (sirm3d):

don't mind the quotient. the result of the operation is the remainder only. FORGET the quotient Q

OpenStudy (freckles):

Now you try 9*9

OpenStudy (freckles):

Remainders shouldn't be negative is why I said take Q to be 1

OpenStudy (freckles):

\[\text{ \check} 10*11=9 =R\] \[\frac{10+11}{12}=\frac{21}{12}=Q+\frac{R}{12}=1+\frac{9}{12}\] 12 goes into 21 1 time

OpenStudy (hba):

Thanks a lot everyone.